# An Iterative Bayesian Approach for System Identification based on Linear Gaussian Models

### v2

@@ IV-D Algorithm
Our algorithm first runs a loop to improve the parameter estimate. We obtain a new parameter estimate by solving problem (8). We only accept the new solution if the linearization errors are insignificant compared to the model errors. In this case, we update the covariance $\Sigma$ and resolve problem (8) using the new covariance and the Taylor expansion around the new parameter estimate. If the linearization errors are deemed significant we reject the new estimate, reduce $\delta$ and resolve. This iterative technique will produce a new parameter estimate. It will further produce an accurate error covariance $\Sigma_{\mathrm{model\ error}}$ for the errors attributed to the discrepancy between the current model and the true system.
@@
Algorithm 1 allows for testing whether the chosen parametric model family $f$ is suitable for the system under consideration. We know that the parameter estimates obtained from the successive calls to algorithm 1 correspond to the best model fits for the obtained datapoints and Taylor MAP objective. Suppose that the magnitude of $\Sigma_{\mathrm{model\ error}}$ does not decrease between successive calls to algorithm 1. This suggests that the chosen parametric family $f$ is not appropriate for the system.
Algorithm 1 Proposed Algorithm for System Identification
@@
4:     Compute $\hat{\epsilon}_{i},\bar{\epsilon}_{i}$ for all $i=1,\dots,n$
5:     if $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ large compared to $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$ then
@@
9:         Set $\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}$
@@ V Results
We test our algorithm on both linear and nonlinear systems. [^1] The code is included in the repository: https://github.com/alextzik/informative_input_design Our results show the importance of finetuning the covariances $\Sigma$ in our algorithm.
Baselines
We consider a variant of our proposed algorithm as baseline. The baseline does not rely on any finetuning. Given the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, $\Sigma$, $\Sigma_{\mathrm{prior}}$ and $\hat{\theta}$, it obtains an updated estimate $\hat{\theta}_{+}$ by solving eq. 8. It then solves eq. 11 to obtain the next input $x_{n+1}$. It does not update $\Sigma$, which is always equal to the initial value provided.
@@
We set $\delta=0.3$. In line 5 of algorithm 1, we reduce $\delta$ by setting $\delta\leftarrow 0.8\delta$. We do not consider any constraint s on $\mathcal{X}$ or $\mathcal{Y}$ for simplicity in the exposition. The first datapoints are assumed fixed for all methods. Further, we run the algorithms with multiple initializations of $\Sigma$. We use the scalar function $\log\det$ as $\mathcal{M}$ when determining the next input.
@@
In most cases below , the true system belongs to the parametric model family. We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number. Further, we can compare the methods with respect to the $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep, which is given in eq. 9. A lower value signifies less uncertainty in the parameter estimate and a more informative sequence of inputs. The results are gathered in Table I for all experiments.
@@ V-A Linear System
Our results are shown in Figure 3. In this experiment, we also bound the norm of the input to be less than $0.5$. In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses. We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, although the complete estimate is 4-dimensional. We further plot the designed inputs for the two methods. Our algorithm exhibits a faster convergence to the true model parameters compared to the baseline. This is verified by the lower value of $\log\det$ for the final error covariance.
@@ V-B Hénon map dynamics
Our results are shown in Figure 4. Our proposed algorithm is able to converge to the true model parameters and outperforms the baseline with respect to both metrics considered. From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter value, the estimate along its trajectory has a higher uncertainty in the first coordinate $\theta_{1}$, which is captured by the elongated ellipse axis along that direction. We also observe that the error ellipses shrink as we move towards the true parameters. This implies that our uncertainty is reduced.
@@ V-C Unicycle Model
$f(x_{t},u_{t};\theta)=\begin{bmatrix}x_{t,1}+u_{1,t}\theta_{1}\cos(x_{t,3})\\ \theta_{2}x_{t,2}+u_{1,t}\theta_{1}\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\theta_{1}\end{bmatrix}.$ (22)
The time discretization is thus one of the unknown parameters. We start the system from the origin with zero being the first control for both methods. Figure 5 shows the performance of our algorithm. It outperforms the considered baseline. The posterior error ellipses are an accurate representation of the uncertainty. Once, $\theta_{1}$ is correctly estimated, the ellipses become vertical, since there is uncertainty only in $\theta_{2}$. We further plot the state trajectory $(x_{t,1},x_{t,2})$ for the two methods. In this experiment, we bound the norm of the input to be less than $1$.
V-D Importance of Finetuning
In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with the baseline. The initial covariance $\Sigma$ is set to $0.2I$, while $\Sigma_{\mathrm{prior}}=0.1I$ and $\theta_{\mathrm{prior}}=(7,4)$. It is evident in Figure 1 that the baseline does not converge properly. This can be attributed to the high confidence we have initially placed on both the prior value of $\theta$ and the accuracy of the measurements. Under these conditions, without finetuning, it is impossible for the algorithm to detect the false confidence. This highlights the importance of finetuning the covariance $\Sigma$.
Figure 1: Importance of finetuning the covariance matrix $\Sigma$ in the case of Hénon dynamics.
V-E Model Mismatch
We consider the case where the parametric model family does not include the true system. We suppose that the true system is the Hénon dynamics of eq. 19 and the parametric model family consists of linear models of the form
@@ V-E Model Mismatch
Obviously, this is an inadequate model family. As discussed in section IV-D, we expect our algorithm to provide an indication of the inadequacy of the parametric model family via the magnitude of $\Sigma_{\mathrm{model\ error}}$. This is shown in Figure 2. The magnitude of $\Sigma_{\mathrm{model\ error}}$, here chosen as $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease as the iterations progress. This implies that the realized errors remain high. Given that our algorithm computes MLEs, this indicates that the obtained parameter estimates, which are in a sense optimal, are inadequate to capture the true dynamics.
Figure 2: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over iterations when the parametric model family cannot capture the true system dynamics. The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease indicates an inadequate parametric model family.
(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
@@
(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
@@
(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
(b) State of dynamics as a function of time for the various methods for a specific initialization of $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
@@
Method/Experiment Linear Dynamics Hénon Dynamics Unicycle Dynamics
proposed $\mathbf{-77.2}$ $\mathbf{-51.2}$ $\mathbf{-43.5}$
baseline $-12.1$ $-31.8$ $-7.5$
TABLE I: $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep for the various dynamics studied.
@@ VI Conclusion
We present a Bayesian methodology for system identification that is based on linear Gaussian assumptions. It extends to general nonlinear system identification through the Taylor expansion. A key module in our algorithm is the online calibration of the noise covariance that greatly impacts performance and allows us to judge the suitability of a parametric model family for the underlying system. Our algorithm also has limitations. The use of the trust region, although necessary to ensure the accuracy of the Taylor series, might forbid large changes in the parameters when these are implied by the data. Future work will apply our method to high-dimensional problems and look into theoretical guarantees.
@@ Acknowledgements
Toyota Research Institute (TRI) provided funds to assist the authors with their research, but this article solely reflects the opinions and conclusions of its authors and not TRI or any other Toyota entity. The NASA University Leadership Initiative (grant ##\##80NSSC20M0163) provided funds to assist the first author with their research, but this article solely reflects the opinions and conclusions of its authors and not any NASA entity. For the first author, this work was also partially funded through the Alexander S. Onassis Foundation Scholarship program. We would like to thank Joshua Ott and Arec Jamgochian for their feedback.

### v3

@@ IV-D Algorithm
Our algorithm first runs a loop to improve the parameter estimate. We obtain a new parameter estimate by solving problem (8). We only accept the new solution if the linearization errors are small compared to the model errors. In this case, we update the covariance $\Sigma$ and re-solve problem (8) using the new covariance and the Taylor expansion around the new parameter estimate. If the linearization errors are deemed significant we reject the new estimate, reduce $\delta$ and re-solve. This iterative technique will produce a new parameter estimate. It will further produce an error covariance $\Sigma_{\mathrm{model\ error}}$ for the errors attributed to the discrepancy between the current model and the true system.
@@
Algorithm 1 allows for testing whether the chosen parametric model family $f$ is suitable for the system under consideration. We know that the parameter estimates obtained from the successive calls to Algorithm 1 correspond to the best model fits for the obtained data-points and the linearized MAP objective. Suppose that the magnitude of $\Sigma_{\mathrm{model\ error}}$ does not decrease between successive calls to Algorithm 1. This suggests that the chosen parametric family $f$ is not appropriate for the system.
Algorithm 1 Proposed Algorithm for Active System Identification
@@
4:  Compute $\epsilon^{\mathrm{model}}_{i},\epsilon^{\mathrm{lin}}_{i}$ for all $i=1,\dots,n$
5:  if $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{lin}}_{i}\rVert$ large compared to $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{model}}_{i}\rVert$ then
@@
9:   Set $\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\epsilon^{\mathrm{model}}_{i}{\epsilon^{\mathrm{model}}_{i}}^{\top}$
@@ V Results
We test our algorithm on both linear and nonlinear systems. [^1] The code is included in the repository: https://github.com/alextzik/informative_input_design Our results show the importance of fine-tuning the covariances $\Sigma$ in our algorithm.
@@
We set $\delta=0.3$. In line 5 of algorithm 1, we reduce $\delta$ by setting $\delta\leftarrow 0.8\delta$. The first data-points are assumed fixed. Furthermore, we independently run the algorithm for $30$ different initializations of $\Sigma$, $\theta_{\mathrm{prior}}$, and $\Sigma_{\mathrm{prior}}$. We use the scalar function $\log\det$ as $\mathcal{M}$ when determining the next input. The loop in lines 1-10 in Algorithm 1 is carried out for $10$ times.
@@
In section V-A, section V-B, and section V-C below , the true system belongs to the parametric model family, i.e., $g(\cdot)=f(\cdot;\theta_{\mathrm{true}})$ for some $\theta_{\mathrm{true}}\in\mathbb{R}^{d_{\theta}}$. We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number. The iteration equals the size of the dataset.
@@ V-A Linear System
Our results are shown in Figure 3. In this experiment, we also bound the norm of the input to be less than $0.5$. Based on the error plot, our algorithm is able to converge to the true parameter values irrespective of the initial $\Sigma$, $\theta_{\mathrm{prior}}$, and $\Sigma_{\mathrm{prior}}$.
In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses for a particular choice of $\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}}$, and initial $\Sigma$. We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, $\hat{\theta}_{1}$ and $\hat{\theta}_{2}$, although the complete estimate is 4-dimensional. We further plot the designed inputs for the same $\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}}$, and initial $\Sigma$.
@@ V-B Hénon map dynamics
Our results are shown in Figure 4. Our proposed algorithm is able to converge to the true model parameters. From the estimate trajectory, we observe that the posterior covariance (9) captures the uncertainty in the parameter: based on the true parameter value, the estimate has larger errors in the first coordinate $\theta_{1}$ at later times, which is captured by the elongated ellipse axis along that direction for those times. We also observe that the error ellipses shrink as we move towards the true parameters. This implies that our uncertainty is reduced.
@@ V-C Unicycle Model
$f(x_{t},u_{t};\theta)=\begin{bmatrix}\theta_{2}x_{t,1}+u_{1,t}\theta_{1}\cos(x_{t,3})\\ \theta_{2}x_{t,2}+u_{1,t}\theta_{1}\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\theta_{1}\end{bmatrix}$. (22)
The time discretization is thus one of the unknown parameters. Figure 5 shows the performance of our algorithm. Once, $\theta_{1}$ is correctly estimated, the ellipses become vertical, since there is uncertainty only in $\theta_{2}$. In this experiment, we bound the norm of the input $u_{1,t}$ to be less than $1$.
V-D Model Mismatch
We consider the case where the parametric model family does not include the true system. We say that the model family does not capture the true dynamics. We suppose that the true system is the Hénon dynamics of eq. 19 and the parametric model family consists of linear models of the form
@@ V-D Model Mismatch
Obviously, this is an inadequate model family. As discussed in section IV-D, we expect our algorithm to provide an indication of the inadequacy of the parametric model family via the magnitude of $\Sigma_{\mathrm{model\ error}}$. This is shown in Figure 1. The magnitude of $\Sigma_{\mathrm{model\ error}}$, here chosen as $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus higher than its initial value. This implies that the realized errors remain high. Lines 4-5 of Algorithm 1 ensure that the large model errors are not because of a bad step in the optimization of (8). In other words, lines 4-5 make it unlikely that we have the appropriate model family, but a bad estimate $\hat{\theta}$ due to the linearization error in problem (8). This indicates that the obtained parameter estimates, which are in a sense optimal, are inadequate to capture the true dynamics. In the case of the model family (20), which captures the dynamics, $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus lower than its initial value. In other words, we suggest the following empirical guideline:
If $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus at a value higher than its initial value and sufficiently large (close to 0) or rises across time, consider modifying the model family. Equivalently, we can look at the model errors $\epsilon_{i}^{\mathrm{model}}$ to draw our conclusions.
We repeat the process for the inadequate model family (18) in Figure 2.
Figure 1: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over time for the parametric model family (23) and the Hénon dynamics. The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus higher than its initial value indicates that this is an inadequate parametric model family.
Figure 2: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over time for the parametric model family (18) and the Hénon dynamics. The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus higher than its initial value indicates that this is an inadequate parametric model family.
(a) Model parameter estimate (point value in red and posterior covariance (9) in blue) as it evolves over time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
(b) Inputs as a function of time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval. The randomness is over $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
@@
(a) Model parameter estimate (point value in red and posterior covariance (9) in blue) as it evolves over time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
(b) Inputs as a function of time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval. The randomness is over $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
@@
(a) Model parameter estimate (point value in red and posterior covariance (9) in blue) as it evolves over time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
(b) Inputs as a function of time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval. The randomness is over $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
@@ VI Conclusion
We present a Bayesian methodology for system identification that is based on linear Gaussian assumptions. It extends to general nonlinear system identification through the Taylor expansion. A key module in our algorithm is the online calibration of the noise covariance that also allows us to judge the suitability of a parametric model family for the underlying system. Our algorithm however has limitations. The use of the trust region, although necessary to ensure the accuracy of the Taylor series, might forbid large changes in the parameters when these are implied by the data. The calibration of the covariance is dependent on the current estimate value, which can slow down performance if the current estimate is not good. Future work will apply our method to high-dimensional problems and look into theoretical guarantees.
@@ Acknowledgements
Toyota Research Institute (TRI) provided funds to assist the authors with their research, but this article solely reflects the opinions and conclusions of its authors and not TRI or any other Toyota entity. The NASA University Leadership Initiative (grant #\#80NSSC20M0163) provided funds to assist the first author with their research, but this article solely reflects the opinions and conclusions of its authors and not any NASA entity. For the first author, this work was also partially funded through the Alexander S. Onassis Foundation Scholarship program. We would like to thank Joshua Ott and Arec Jamgochian for their feedback.